# Enhanced Pruning Strategy for Multi-Component Neural Architectures Using Component-Aware Graph Analysis

### v1

@@ Abstract
Deep neural networks (DNNs) deliver outstanding performance, but their complexity often prohibits deployment in resource-constrained settings. Comprehensive structured pruning frameworks based on parameter dependency analysis reduce model size with specific regard to computational performance. When applying them to Multi-Component Neural Architectures (MCNAs), they risk network integrity by removing large parameter groups. We introduce a component-aware pruning strategy, extending dependency graphs to isolate individual components and inter-component flows. This creates smaller, targeted pruning groups that conserve functional integrity. Demonstrated effectively on a control task, our approach achieves greater sparsity and reduced performance degradation, opening a path for optimizing complex, multi-component DNNs efficiently.
@@
While the tremendous advance of deep learning yields increasingly complex and high-performing neural networks, especially for visual data, a critical conflict emerges. The very attributes driving this success—large model sizes, demanding computations, and intensive matrix operations—clash directly with the limitations of consumer electronics and edge devices such as microcontrollers, hindering practical deployment where resources are scarce.
@@
In architectures where different network modules serve different purposes, e.g., encoders, prediction models, and control policies as in TD-MPC (Hansen et al., 2022) or JEPA (Assran et al., 2023; Ulmen et al., 2025), dependency graphs, which are commonly used for pruning, cannot capture the intricate data flows between the components. Unlike monolithic networks, such Multi-Component Neural Architectures (MCNA) demand a more granular dependency model to represent the intra- and inter-component interactions accurately. Additionally, although all components of MCNAs have to be trained, not all components are used in downstream tasks, which should be accounted for when prioritizing which component to prune.
@@ 2 Related Work
Simple magnitude-based pruning, which removes weights with the smallest absolute values, was one of the earliest techniques. Despite its simplicity, this approach has shown surprising effectiveness, often matching or outperforming more complex methods (Gupta et al., 2022). However, the field has expanded to include structured pruning. Structured pruning removes entire sets of parameters, enabling direct computational and memory savings. This approach has become particularly important for deploying models on edge devices (He and Xiao, 2023). Structurally Prune Anything (SPA) has been developed to prune any architecture from any framework and at any stage of training (Wang et al., 2024).
Recent advancements in pruning offer more sophisticated parameter evaluation through Hessian-based and class-aware methods (Chong et al., 2023; Jiang et al., 2023). To counteract potential accuracy degradation from pruning, techniques like knowledge distillation are employed to boost performance (Qian et al., 2024). Concurrently, sparse optimization techniques integrate regularization during pretraining and pruning, achieving competitive results without needing subsequent fine-tuning (Shi et al., 2024). Theoretical progress is also evident, with operator-theoretic perspectives unifying magnitude and gradient-based pruning (Redman et al., 2022). Pruning is further integrated with gradient information (Yang et al., 2024) and quantization, enabling automatic trade-offs for hardware efficiency (Wang et al., 2020). Finally, reflecting the growing emphasis on sustainable AI (Cheng et al., 2024), energy-constrained optimization methods explicitly target energy reduction by formulating pruning under energy budget constraints (Liu et al., 2022). These methods - and to our knowledge, no existing method - try to formalize pruning concerning the functional and structural interdependence of different network components.
Structured pruning’s hardware compatibility facilitates real-time inference, aided by compiler optimizations and fine-grained methods (Gong et al., 2021). Frameworks like HALP globally optimize accuracy under latency constraints (Shen et al., 2022). Demonstrated speedups include 2x for YOLOv7 (FPGA) (Pavlitska et al., 2024), 19.1x for LSTM (Jetson) (Lindmar et al., 2022-06-13), and 2x for GNNs (PruneGNN on A100) (Gurevin et al., 2024). It enhances compute-in-memory (11.1x compression) (Meng et al., 2022) and systolic arrays (SCRA, 4.79x speedup) (Zhang et al., 2023). Hardware-architecture co-design approaches like CSP further improve efficiency by preserving data reuse and utilizing predictable sparsity (Hanson et al., 2022).
@@ 3.1 Model Compression for Neural Network Controllers: Challenges Beyond Standard MCNAs
In this study, we specifically reference the latter and more stringent scenario to highlight the need for more intricate implications.
In the latter case, even minimal component-agnostic pruning can disrupt system behavior so that the imposed requirements are no longer met. A fundamental relationship governing component interaction is their functional sequence within the network’s execution flow, where each step relies on the sufficient performance of its predecessor. For example, in the aforementioned architectures TD-MPC and JEPA, a slight alteration of the encoder’s parameters can severely deform the learned manifold and adversely affect the succeeding world model by providing an unfamiliar latent representation. This interdependency underscores that downstream module fidelity critically depends on the output quality of upstream components. Thus, effective compression must extend beyond mere structural simplification.
@@ 3.3 Further Challenges in MCNA
In complex MCNAs, module interactions extend far beyond simple feedforward links, often incorporating feedback loops, recurrent connections, and dynamic gating. Standard DepGraph, which relies on static, direct dependencies, oversimplifies these relationships. Feedback loops, for example, can either obscure key dependencies or cause multiple layers to merge into oversized groups. Recurrent connections—such as planning modules that repeatedly invoke a dynamics model—might not be properly captured, misaligning the temporal dependencies critical for proper function. Likewise, gating and conditional execution introduce dynamic connectivity that DepGraph fails to distinguish. These limitations lead DepGraph to treat all components equally, neglecting their unique functionalities, sensitivities and interdependencies, and ultimately compromising pruning effectiveness.
@@ 4.3 Inter-Component Interface Dependency ()
This term addresses dependencies between layers residing in different components (i.e., ). It is activated when the layer produces the designated output interface of component and layer consumes the input interface of component . There exists a verified connection between the interfaces ().
@@ 4.5 Identifying Components and Their Dependencies
Dynamic tracing is performed using PyTorch forward and backward hooks. The forward hooks record the runtime inputs and outputs along their dimensions of individual components, while backward hooks captures gradient flows to make sure that the structural integrity is not lost and further training is still possible. The forward and backward flows are the basis for adjustments to the dependency edges.
The dependency graph obtained can be further verified through certain techniques. Shape matching ensures that the output dimension from a component must match the corresponding receiver’s input dimension, and if multiple components contribute to an input, their combined output dimensions should equal the total input dimension of the receiving component. This rules out any dimensional discrepancies that suggest missing or erroneous dependencies. Dynamic tensor tracing via the PyTorch autograd graph records the lineage of intermediate tensors during each forward pass. In cases of non-standard data manipulations (e.g., reshaping, slicing, aggregating) or when components share buffers or weights, semantic heuristics instantiate plausible connections again based on the network’s intended functionality.
@@ 4.6 Pruning Groups and Importance Calculation
Isolating pruning groups on a per-component basis confines the scope of pruning decisions, preventing an infeasible change of the dimensions between connected components. Additionally, we enable the expert user to exclude certain components from (further) pruning, e.g., to prevent a further removal of parameters in an upstream component, so that performance degradation does not cascade to downstream components. For each group, we compute an importance metric, utilizing or norms of the weights to determine the order of pruning.
@@ 5 Experiments and Discussion
In the second experiment, we investigated our grouping approach by comparing the vanilla dependency graph method with our component-awareness. The vanilla method produced fewer total groups with larger average group sizes, indicating coarser groupings. In contrast, our method generated a higher number of groups with smaller sizes; for instance, in the Branched model, the vanilla method resulted in 20 groups with an average group size of 4.10 operations, whereas our component-aware method produced 33 groups with an average size of 1.73 operations. This demonstrates that our approach yields finer, more precise partitions of the architecture.
To further assess whether creating finer groups translates into tangible benefits in both performance preservation and model size reduction, we conducted a final experiment using a TD-MPC model on the cartpole-swingup Gym environment with image inputs (Brockman et al., 2016). An over-parameterized TD-MPC model (with each component having a width of 512 and a depth of three layers) was first trained until the average reward stabilized around 850. We then applied pruning to reduce the amount of parameters by a range of 5% to 80%, using both the baseline (vanilla dependency graph) and our component-aware method. As shown in Figure 2, while increasingly pruning parameters naturally decreases the rewards, the vanilla approach experienced an immediate decrease and an abrupt 50% drop at 40% of parameter reduction. In contrast, our component-aware technique maintained a gradual decline, hinting at its robustness.
Moreover, by revealing the constituent components within each pruning group, our method offers the flexibility to protect sensitive components from being overly pruned. This was demonstrated by assigning higher importance weighting to groups associated with the encoder in the TD-MPC model, thereby reducing the likelihood of excessive pruning in these critical regions. Overall, these results substantiate that our component-aware approach not only provides a more granular grouping of the MCNA architectures but also preserves performance more effectively, even during aggressive pruning. This marks a significant advancement over existing methods by enabling model size reduction without a substantial loss in performance.
@@ 6 Conclusion
In this work, we demonstrated that switching to a component-aware dependency graph produces smaller, targeted, structured pruning groups while conserving functional integrity. Our method, which identifies network components and accurately models both intra- and inter-dependencies, leads to a reduced average group size and thereby mitigates the overall pruning impact. Because we ran all inference on a standard PC in PyTorch’s eager mode, we observed only modest speed‑ups as pruning increased; in future work, we will explore inference‑time gains using embedded hardware and optimized inference engines.
Our ongoing research is focused on expanding the importance criteria of these groups by incorporating sensitivity analyses and theoretical constraints, including stability guarantees. Future work will explore the benefits of post‑pruning fine‑tuning to further enhance performance, as well as the integration of component‑aware pruning with other model reduction techniques, such as quantization, to reduce model size without compromising performance.

### v2

@@ Abstract
Deep neural networks (DNNs) deliver outstanding performance, but their complexity often prohibits deployment in resource-constrained settings. Comprehensive, structured pruning frameworks based on parameter dependency analysis reduce model size with specific regard to computational performance. When applying them to Multi-Component Neural Architectures (MCNAs), they risk network integrity by removing large parameter groups. We introduce a component-aware pruning strategy, extending dependency graphs to isolate individual components and inter-component flows. This creates smaller, targeted pruning groups that conserve functional integrity. Demonstrated effectively on a control task, our approach achieves greater sparsity and reduced performance degradation, opening a path for efficiently optimizing complex, multi-component DNNs.
@@
While deep learning’s tremendous advances yield increasingly complex and high-performing neural networks, especially for visual data, a critical conflict emerges. The very attributes driving this success—large model sizes, demanding computations, and intensive matrix operations—clash directly with the limitations of consumer electronics and edge devices such as microcontrollers, hindering practical deployment where resources are scarce.
@@
In architectures where different network modules serve different purposes, e.g., encoders, prediction models, and control policies, as in TD-MPC (Hansen et al., 2022) or JEPA (Assran et al., 2023; Ulmen et al., 2025), dependency graphs, which are commonly used for pruning, cannot capture the intricate data flows between the components. Unlike monolithic networks, such Multi-Component Neural Architectures (MCNA) demand a more granular dependency model to represent the intra- and inter-component interactions accurately. Additionally, although all components of MCNAs have to be trained, not all components are used in downstream tasks, which should be accounted for when prioritizing which component to prune.
@@ 2 Related Work
Simple magnitude-based pruning, which removes weights with the smallest absolute values, was one of the earliest techniques. Despite its simplicity, this approach has shown surprising effectiveness, often matching or outperforming more complex methods (Gupta et al., 2024). However, the field has expanded to include structured pruning. Structured pruning removes entire sets of parameters, enabling direct computational and memory savings. This approach has become particularly important for deploying models on edge devices (He and Xiao, 2023). Structurally Prune Anything (SPA) has been developed to prune any architecture from any framework and at any stage of training (Wang et al., 2023).
Recent advancements in pruning offer more sophisticated parameter evaluation through Hessian-based and class-aware methods (Chong et al., 2023; Jiang et al., 2024). To counteract potential accuracy degradation from pruning, techniques like knowledge distillation are employed to boost performance (Qian et al., 2024). Concurrently, sparse optimization techniques integrate regularization during pretraining and pruning, achieving competitive results without needing subsequent fine-tuning (Shi et al., 2024). Theoretical progress is also evident, with operator-theoretic perspectives unifying magnitude and gradient-based pruning (Redman et al., 2022). Pruning is further integrated with gradient information (Yang et al., 2024) and quantization, enabling automatic trade-offs for hardware efficiency (Wang et al., 2020). Finally, reflecting the growing emphasis on sustainable AI (Cheng et al., 2024), energy-constrained optimization methods explicitly target energy reduction by formulating pruning under energy budget constraints (Liu et al., 2022). These methods - and to our knowledge, no existing method - try to formalize pruning concerning the functional and structural interdependence of different network components.
Structured pruning’s hardware compatibility facilitates real-time inference, aided by compiler optimizations and fine-grained methods (Gong et al., 2022). Frameworks like HALP globally optimize accuracy under latency constraints (Shen et al., 2022). Demonstrated speedups include 2x for YOLOv7 (FPGA) (Pavlitska et al., 2024), 19.1x for LSTM (Jetson) (Lindmar et al., 2022), and 2x for GNNs (PruneGNN on A100) (Gurevin et al., 2024). It enhances compute-in-memory (11.1x compression) (Meng et al., 2022) and systolic arrays (SCRA, 4.79x speedup) (Zhang et al., 2023). Hardware-architecture co-design approaches like CSP further improve efficiency by preserving data reuse and utilizing predictable sparsity (Hanson et al., 2022).
@@ 3.1 Model Compression for Neural Network Controllers: Challenges Beyond Standard MCNAs
In the latter case, even minimal component-agnostic pruning can disrupt system behavior so that the imposed requirements are no longer met. A fundamental relationship governing component interaction is their functional sequence within the network’s execution flow, where each step relies on the sufficient performance of its predecessor. For example, in the aforementioned architectures, TD-MPC and JEPA, a slight alteration of the encoder’s parameters can severely deform the learned manifold and adversely affect the succeeding world model by providing an unfamiliar latent representation. This interdependency underscores that downstream module fidelity critically depends on the output quality of upstream components. Thus, effective compression must extend beyond mere structural simplification.
@@ 3.3 Further Challenges in MCNA
In complex MCNAs, module interactions extend far beyond simple feedforward links, often incorporating feedback loops, recurrent connections, and dynamic gating. Standard DepGraph, which relies on static, direct dependencies, oversimplifies these relationships. Feedback loops, for example, can either obscure key dependencies or cause multiple layers to merge into oversized groups. Recurrent connections—such as planning modules that repeatedly invoke a dynamics model—might not be properly captured, misaligning the temporal dependencies critical for proper function. Likewise, gating and conditional execution introduce dynamic connectivity that DepGraph fails to distinguish. These limitations lead DepGraph to treat all components equally, neglecting their unique functionalities, sensitivities, and interdependencies, and ultimately compromising pruning effectiveness.
@@ 4.3 Inter-Component Interface Dependency ()
This term addresses dependencies between layers residing in different components (i.e., ). It is activated when layer produces the designated output interface of component and layer consumes the input interface of component . There exists a verified connection between the interfaces ().
@@ 4.5 Identifying Components and Their Dependencies
Dynamic tracing is performed using PyTorch forward and backward hooks. The forward hooks record the runtime inputs and outputs along their dimensions of individual components, while backward hooks capture gradient flows to make sure that the structural integrity is not lost and further training is still possible. The forward and backward flows are the basis for adjustments to the dependency edges.
The dependency graph obtained can be further verified through certain techniques. Shape matching ensures that the output dimension from a component must match the corresponding receiver’s input dimension, and if multiple components contribute to an input, their combined output dimensions should equal the total input dimension of the receiving component. These rule out any dimensional discrepancies that suggest missing or erroneous dependencies. Dynamic tensor tracing via the PyTorch autograd graph records the lineage of intermediate tensors during each forward pass. In cases of non-standard data manipulations (e.g., reshaping, slicing, aggregating) or when components share buffers or weights, semantic heuristics instantiate plausible connections again based on the network’s intended functionality.
@@ 4.6 Pruning Groups and Importance Calculation
Isolating pruning groups on a per-component basis confines the scope of pruning decisions, preventing an infeasible change of the dimensions between connected components. Additionally, we enable the expert user to exclude certain components from (further) pruning, e.g., to prevent further removal of parameters in an upstream component, so that performance degradation does not cascade to downstream components. For each group, we compute an importance metric, utilizing or norms of the weights to determine the order of pruning.
@@ 5 Experiments and Discussion
In the second experiment, we investigated our grouping approach by comparing the vanilla dependency graph method with our component awareness. The vanilla method produced fewer total groups with larger average group sizes, indicating coarser groupings. In contrast, our method generated a higher number of groups with smaller sizes; for instance, in the Branched model, the vanilla method resulted in 20 groups with an average group size of 4.10 operations, whereas our component-aware method produced 33 groups with an average size of 1.73 operations. This demonstrates that our approach yields finer, more precise partitions of the architecture.
To further assess whether creating finer groups translates into tangible benefits in both performance preservation and model size reduction, we conducted a final experiment using a TD-MPC model on the cartpole-swingup Gym environment with image inputs (Brockman et al., 2016). An over-parameterized TD-MPC model (with each component having a width of 512 and a depth of three layers) was first trained until the average reward stabilized around 850. To evaluate its effectiveness, we compared our component-aware method against the baseline (vanilla dependency graph) approach across various sparsity levels, from 5% to 80%. The results, presented in Figure 2, reveal a sheer difference in performance degradation. The baseline method suffered an immediate performance drop, culminating in a catastrophic 50% reward loss at just 40% sparsity. In contrast, our component-aware approach exhibited a much more graceful and gradual decline, demonstrating its superior robustness.
Moreover, by revealing the constituent components within each pruning group, our method offers the flexibility to protect sensitive components from being overly pruned. This was demonstrated by assigning higher importance weighting to groups associated with the encoder in the TD-MPC model, thereby reducing the likelihood of excessive pruning in these critical regions. Overall, these results substantiate that our component-aware approach not only provides a more granular grouping of the MCNA architectures but also preserves performance more effectively, even during aggressive pruning. This represents a significant advancement over existing methods, enabling model size reduction without a substantial loss in performance.
@@ 6 Conclusion
In this work, we demonstrated that switching to a component-aware dependency graph produces smaller, targeted, structured pruning groups while conserving functional integrity. Our method, which identifies network components and accurately models both intra- and inter-dependencies, leads to reduced average group size and thereby mitigates the overall pruning impact. Because we ran all inference on a standard PC in PyTorch’s eager mode, we observed only modest speed‑ups as pruning increased; in future work, we will explore inference‑time gains using embedded hardware and optimized inference engines.
Our ongoing research is focused on expanding the importance criteria of these groups by incorporating sensitivity analyses and theoretical constraints, including stability guarantees. Future work will explore post-pruning fine-tuning to enhance performance further, as well as the integration of component-aware pruning with other model reduction techniques like quantization. To conclusively demonstrate the method’s impact and scalability, we will then broaden our trials to more complex architectures, such as Large Language Models (LLMs), and a diverse range of applications.